A box of books weighing 316 N moves with
a constant velocity across the floor when it is
pushed with a force of 376 N exerted downward at an angle of 31.2◦ below the horizontal.
Find µk between the box and the floor.
1
Expert's answer
2019-12-24T14:30:32-0500


From the above

Horizontal component of the force will be "F\cos\theta= 376 \cos31.2^\circ"

Vertical component ="376\sin 31.2\circ+mg=376\sin32.2+316"

"=194.78+316=510.78N"

As per the condition given the question, that it is moving with the constant velocity, it means that the net resultant force is zero.

Hence,

"376\cos 31.2^o=\mu N"

"\mu=\dfrac{376\cos31.2^o}{510.78}=\dfrac{321.6}{510.78}=0.63"